OpenStudy (anonymous):

A fossilized animal that was killed by a volcano eruption has 12.5 percent of its original amount of carbon-14 remaining; this radioisotope has a half-life of 5730 years. How long ago did the volcano erupt? 34,380 years ago 22,920 years ago 17,190 years ago 9,168 years agoA fossilized animal that was killed by a volcano eruption has 12.5 percent of its original amount of carbon-14 remaining; this radioisotope has a half-life of 5730 years. How long ago did the volcano erupt? 34,380 years ago 22,920 years ago 17,190 years ago 9,168 years ago

jimthompson5910 (jim_thompson5910):

is this the same question? just posted twice?

OpenStudy (anonymous):

Yes sorry! I posted it in the wrong category at first

jimthompson5910 (jim_thompson5910):

that's fine, just curious

jimthompson5910 (jim_thompson5910):

use the formula F = P(1/2)^(t/H) where F = final amount in t years P = initial amount t = time in years H = halflife

jimthompson5910 (jim_thompson5910):

in this case F = 0.125 P = 1 t = unknown (this is what we want to solve for) H = 5370

OpenStudy (anonymous):

Im so confused! I just dont understand this lesson at all and I was sick so missed the lesson in class.

jimthompson5910 (jim_thompson5910):

this looks completely foreign?

OpenStudy (anonymous):

Ok i got 16110? am i close?

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

oh sry meant to say H = 5730

OpenStudy (anonymous):

ok so 17190 would be correct? what about the other 2 questions>

jimthompson5910 (jim_thompson5910):

F = P(1/2)^(t/H) 0.125 = 1*(1/2)^(t/5730) 0.125 = (1/2)^(t/5730) ln(0.125) = ln((1/2)^(t/5730)) ln(0.125) = (t/5730)*ln(1/2) 5730*ln(0.125)/ln(1/2) = t t = 5730*ln(0.125)/ln(1/2) t = 17,190

jimthompson5910 (jim_thompson5910):

yep it's 17190

OpenStudy (anonymous):

thank you so much I figured them out!

jimthompson5910 (jim_thompson5910):

you're welcome